OpenStudy (anonymous):

Hi, I want to how to indefinite (8u^3-1)/(2u/1)?? Thank you for answering........

OpenStudy (id21):

check whether you typed expression is correct. really (2u/1)? because it is the sames as 2u.

OpenStudy (anonymous):

sorry a typical mistake, it's 2u-1...... the question is (8u^3-1)/(2u-1)....... thank you for answering it

OpenStudy (id21):

Use formula \[a^3-b^3=(a-b)(a^2+ab+b^2)\] \[\frac{2^3u^3-1^3}{2u-1}=\frac{(2u-1)(4u^2+2u+1)}{2u-1}=4u^2+2u+1\]

OpenStudy (anonymous):

thank you